Write an equation in slope intercept form of the line having the given slope and passing through the given point: m=0,(0,3)
This is my question..

Answer by stanbon(57361)

Write an equation in slope intercept form of the line having the given slope and passing through the given point: m=0,(0,3)
------------
y=mx+b
You are told m=0 and b=3
EQUATION:
y=0(x)+3
y=3
Cheers,
Stan H.